No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>

Reminder

To have a better experience, please upgrade your IE browser.

upgrade

Improper setting of optical module temperature alarm threshold makes the T2SL64 board report the TEM_HA alarm

Publication Date:  2012-07-24 Views:  34 Downloads:  0
Issue Description
In the operation of the network of customer A in country G, the SST2SL64 board of several OSN7500 sites reports the TMP_HA alarm and the service runs normally      
Alarm Information
TEM_HA      
Handling Process
 
Run the following OPTP command:
:optp:1c,0,9b,1,08,13,1,4
Optp cmd : 0813 
01 04 00 e7 00 bf 
Total records :1
In the foregoing commands, “1c” means the slot is 28, the returned value “00e7” indicates the upper threshold, and “00bf” indicates the lower threshold. We find that the set threshold is 19.1~23.1°C. However, the actual performance value is 23.7°C. It indicates that the alarm reporting is normal, and the threshold value of the optical module is set improperly. As confirmed, the site cannot load the threshold value of modifying the optical module. The TEM_HA alarm can be shielded to solve the problem temporarily. To solve the problem thoroughly, the board needs to be replaced.       
Root Cause
The TEM_HA alarm indicates too high temperature of laser. We query the board performance, and the tube core temperature performance (ospitmpcur) value of the board module is 23.7. The service runs normally. The temperature should fall within the working temperature range. We suspect that the temperature control circuit inside the optical module of the board is faulty, or the threshold value is improper, and leads to abnormal reporting of temperature performance alarm of the module.      
Suggestions
None      

END